A bullet (m = 0.0290 kg) is fired with a speed of 99.00 m/s and hits a block (M = 2.40 kg) supported by two light strings as shown, stopping quickly.

A bullet (m = 0.0290 kg) is fired with a speed of


(a) Find the height to which the block rises.

(b) Find the angle (in degrees) through which the block rises, if the strings are 0.330 m in length.
Answer #1

1)use conservation of momentum to get velocity after the collision

(.029)(99)=(2.429)*v
v=1.182 m/s

Now we use conservation of energy to get the height

KE=PE
(1/2)mv2=mgh

h=v2/(2g)

h=1.1822/(2*9.8)

h=.07128 m

2)

h=L(1-cos?)

cos?=(L-h)/(L)

cos?=(.330-.07128)/.330

?=38.37 degrees

Hope that helps